OpenStudy (prominer10000):

Which ordered pair is the solution to the system of equations? x+3y=12 -4x+3y=-3 A. (0, 4) B. (6, 2) C. (3, 3) D. (0, −1)

OpenStudy (eliesaab):

x+3y=12 -4x+3y=-3 x+3y=12 4x-3y=3 5 x = 15 x=3

OpenStudy (eliesaab):

What is the answer?

OpenStudy (prominer10000):

Is it C?

OpenStudy (eliesaab):

Yes

OpenStudy (eliesaab):

Are you understanding what I am doing?

OpenStudy (prominer10000):

I am. :) Thanks for helping! This was my last question.

OpenStudy (eliesaab):

NP